# General Exploratory Bonus for Optimistic Exploration in RLHF

### v1

@@ Acknowledgement
The authors would like to thank Froilan Choi and Shawn Im for their valuable comments on the manuscript. This work is supported in part by the AFOSR Young Investigator Program under award number FA9550-23-1-0184, National Science Foundation under awards IIS-2237037 and IIS-2331669, Office of Naval Research under grant number N00014-23-1-2643, Schmidt Sciences Foundation, Open Philanthropy, Alfred P. Sloan Fellowship, and gifts from Google and Amazon.
@@
We have included all implementation details, hyperparameters, and training procedures in the paper and appendix. Our code and scripts for reproducing the experiments are publicly available through the GitHub repository.
@@
This work studies reinforcement learning from human feedback (RLHF) using only publicly available or synthetic data, without new human subject collection. Here, by providing a rigorous theoretical framework with strong empirical evidence, we pursue a high standard of scientific excellence. We also take into account inclusiveness to make all our visualizations accessible to the unprivileged group of people, by producing figures distinguished by light, shade, and marker. While RLHF has the potential to amplify biases or harmful behaviors if misused, our work is intended solely to advance safe and responsible research, and we encourage its application in alignment with ethical standards.
@@ Appendix A Related Works
Alignment (Hendrycks et al., 2021; Leike et al., 2018; Zhang et al., 2024b; 2025; Yeh et al., 2025) aims to ensure AI systems act in accordance with human values, preferences, and goals; and it has become a critical field in AI research. To steer language models to match human preferences, reinforcement Learning from Human Feedback (RLHF) (Stiennon et al., 2020; Ouyang et al., 2022) acheives great success and has become the standard alignment pipeline. However, its computational complexity has motivated a family of Direct Preference Optimization (DPO) (Rafailov et al., 2024; Azar et al., 2024; Ethayarajh et al., 2024) that forgo explicit reward modeling. Despite their efficiency, recent researchers (Tajwar et al., 2024; Xu et al., 2024; Xiong et al., 2024b) reemphasize the significance of online sampling.
@@
To address the computational overheads of passive exploration in RLHF, which samples trajectories just based on randomness, some existing attempts have been devoted to sample-efficient RL algorithms. Most of works (Das et al., 2024; Ji et al., 2024; Dwaracherla et al., 2024; Mehta et al., 2023; Muldrew et al., 2024) adhere to the principle of optimism, proposing specialized prompt or response selection strategies to emphasize uncertain samples. While some research (Liu et al., 2024; Lou et al., 2024) propose uncertainty-aware reward models with epistemic neural networks or bootstrap ensembles, these methods introduce additional cost. Some research also addresses the sample efficiency with different theoretical foundations, such as information theory (Qi et al., 2025), preference-incentive exploration (Chen et al., 2025). Notably, several works (Zhang et al., 2024a; Xie et al., 2024; Cen et al., 2025) introduce different exploratory bonuses, which can implement optimism toward uncertainty without additional computes. However, they only focus on KL-divergence and their theoretical framework cannot result in real optimism as shown in Section 3.2.
@@
Beyond optimistic exploration, some research proposes fine-grained signals for RL learning. For instance, several research propose segment-level (Kong et al., 2025; Yin et al., 2025) or token-level (Zeng et al., 2024; Li et al., 2024) reward function for alignment or text control. Notably, for reasoning tasks, process reward model (Lightman et al., 2024; Li & Li, 2025; Yuan et al., 2024) which provides step-wise feedback for solutions has shown promise effectiveness. On the other hand, recent research (Cheng et al., 2025; Zheng et al., 2025; Wu et al., 2025) on LLM reasoning reveal that high-entropy tokens guide the model toward diverse reasoning paths. Training with only high-entropy tokens are more beneficial for reasoning performance (Wang et al., 2025). While our approach is highly extensible, we believe the orthogonal methods can be further incorporated with our general exploratory bonus.
@@ Appendix B Optimism Failure of previous works
B.1 Extension beyond $\alpha$-divergence
@@
When $f$ belongs to $f$-divergence, and the reward function is obtained by $\hat{r}=\arg\min_{r}[\mathcal{L}_{BT}(\mathcal{D}_{t},r)-\kappa\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r)]$ and the policy is updated by $\arg\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r_{t})$, the bonus term $\kappa\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r)$ induces the policy model $\pi$ to coincide with $\pi_{\textrm{ref}}$ when $xf^{\prime\prime}(x)$ is a monotone function.
Proof  By Lemma 3.2, we reparameterize the bonus term for optimistic reward-modeling to Eq. 8. Denote $h(u)=uf^{\prime}(u)-f(u)$. For a fixed prompt $x$, we formulate the training process as a constrained problem as follows,
@@
where $\mu,\eta$ are the dual variables. Then we utilize the Karush-Kuhn-Tucker (KKT) conditions for the given optimization problem. The complementary slackness requires that $\forall y,\eta(y)\pi(y|x)=0$. The stationary condition requires
@@
We denote $S_{y}=\{y|\pi(y|x)>0\}$, we have $\forall y\in S_{y},\eta(y)=0$. Since $h^{\prime}(u)=uf^{\prime\prime}(u)>0$ due to the convexity of $f(\cdot)$, we can obtain $\forall y\in S_{y},\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)}$ is a constant. Then applying the normalisation constraint, $\mathbb{E}_{y\sim\pi_{\textrm{ref}}(\cdot|x)}\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)}=1$. Hence, the unique interior optimum is $\pi^{*}(y|x)=\pi_{\textrm{ref}}(y|x)$. $\square$
@@
B.2 Prior adaptions & approximations cannot generalize
Though the theoretical framework of the prior exploratory bonus fails, their empirically implemented loss remains effective through different adaptations and approximations. However, in this subsection, we show that these adaptations and approximations are inextensible beyond the KL-divergence class.
Zhang et al. (2024a) adapt the formulation of $\mathcal{J}_{\beta,f}(\pi,r)$ in $\kappa\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r)$ as
@@ B.2 Prior adaptions & approximations cannot generalize
which adds a bias in the reward expectation term. Under KL-divergence, the original $\mathcal{J}_{\beta,f}(\pi,r)$ will be zero after re-parameterization as shown in Lemma 3.1, thus the sole reparameterized bias term will remain as $\mathcal{E}_{y^{\prime}\sim\pi_{\textrm{ref}}}\log\pi(y^{\prime}|x)$. Since $\mathcal{J}_{\beta,f}(\pi,r)$ cannot be reparameterized to zero except for the KL-divergence, this adaptation cannot generalize then.
In the derivations of Cen et al. (2025), it utilizes an ideal distribution $\pi_{cal}$ which should satisfy $\mathbb{E}_{y\sim\pi_{cal}}r(x,y)=0$. Since $\pi_{cal}$ is practically unobtainable, it uses the rejected responses to approximate $\mathbb{E}_{\pi_{cal}}$, which does not satisfy the predefined condition of $\pi_{cal}$, thus not rigorously coherent to the theory. While the regret decomposition of Xie et al. (2024) relies on the logarithm form of the KL-divergence, it is thus inextensible to a broader divergence class.
@@
In contrast, our general exploratory bonus can seamlessly incorporate an iterative online RLHF algorithm and can naturally extend to the entire $\alpha$-divergence class. All prior bonuses mentioned above can be encompassed by our theoretical framework.
Appendix C Regret Bound
In derivations, we utilize the theoretical tools in Zhang et al. (2024a); Xie et al. (2024); Cen et al. (2025). First, we make some standard statistical assumptions following Cen et al. (2025).
Assumption C.1
For a reward function $r$, and a random function $R(\cdot)$, and any trajectory $\tau$, we have $R_{max}<r(\tau),R(r(\tau))<R_{max}$, where $R_{max}$ is a constant.
This is an assumption generally made for theoretical analyses of RLHF. Note that $R_{max}$ is measurable and controllable in practice. Then we introduce the assumption of the reward class proposed in Cen et al. (2025), which offers a regularization mechanism to incorporate additional policy preferences in the subsequent derivations.
Assumption C.2
We assume that $r^{*}\in\mathcal{R}$, where
$\mathcal{R}=\{r:\mathbb{E}_{x\sim\rho,\tau\sim\pi_{cal}(\cdot|x)}r(x,y)=0\},$ (17)
where $\rho$ is the prompt distribution and $\pi_{cal}$ is a fixed calibration distribution independent of the algorithm.
We also introduce the preference generalized eluder coefficient proposed in Zhang et al. (2024a), an extension of the generalized eluder coefficient (Zhong et al., 2022), which connects prediction error and in-sample estimation error.
Definition C.1
Let $f_{r}(x,y,y^{\prime})=r_{t}(x,y)-r^{*}(x,y)$. For a reward function class $\mathcal{R}$, we define the preference generalized eluder coefficient as the smallest $d_{\mathrm{PGEC}}$ as
$\sum_{t=1}^{T}$ $\mathbb{E}_{x\sim\rho,y\sim\pi^{t}(\cdot|x),y^{\prime}\sim\pi_{cal}}[f_{r_{t}}(x,y,y^{\prime})-f_{r^{*}}(x,y,y^{\prime})]$
$\leq\sqrt{d_{\mathrm{PGEC}}\sum_{t=1}^{T}\mathbb{E}_{x\sim\rho,y\sim\widetilde{\pi}_{t}(\cdot|x),y^{\prime}\sim\pi_{cal}}[f_{r_{t}}(x,y,y^{\prime})-f_{r^{*}}(x,y,y^{\prime})]^{2}}+4\sqrt{d_{\mathrm{PGEC}}T},$ (18)
With the above assumptions and the theoretical tool, we can have the following regret boundary.
Theorem C.1
Let $\mathcal{J}_{\beta,f}(\pi,r)=\mathbb{E}_{x\sim\rho,y\sim\pi(\cdot|x)}r(x,y)-\beta\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}(\cdot|x)}f(\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)})$, when the hyperparameter of the loss Eq. 3 $\kappa=\sqrt{\frac{\log(T|\mathcal{R}|\delta^{-1})}{(\gamma d_{\mathrm{PGEC}}T)}}(32R_{max}e^{4R_{max}})^{-1}$, with probability at least $1-\delta$, the regret can be bounded as follows,
$\sum_{t=1}^{T}\mathcal{J}_{\beta,f}(\pi^{*},r^{*})-\mathcal{J}_{\beta,f}(\pi^{t},r^{*})\leq\mathcal{O}(R_{max}e^{4R_{max}}T\sqrt{d_{\mathrm{PGEC}}\gamma\log(|\mathcal{R}|\delta^{-1})}),$ (19)
where $\gamma=\sup_{x,y}\frac{\pi}{\pi_{cal}}$, $r^{*}$ and $\pi^{*}$ are ground-truth reward function and corresponding optimal policy with $\pi^{*}=\arg\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r^{*})$.
Proof  First, we can decompose the regret function as in Cen et al. (2025) as follows,
$\underbrace{\sum_{t=1}^{T}[\mathcal{J}_{\beta,f}(\pi^{*},r^{*})-\mathcal{J}_{\beta,f}(\pi^{t},r^{t})]}_{\textrm{Term 1}}+\underbrace{\sum_{t=1}^{T}[\mathcal{J}_{\beta,f}(\pi^{t},r^{t})-\mathcal{J}_{\beta,f}(\pi^{t},r^{*})]}_{\textrm{Term 2}}.$ (20)
Then, we will bound term 1 and term 2 individually and combine them at last.
Bound term 1.
First, we connect the term 1 with $\max_{\pi}\mathcal{J}_{\beta,f}(\pi,R(r_{t}))-\max_{\pi}\mathcal{J}_{\beta,f}(\pi,R(r^{*}))$. When $\pi^{*}$ is the optimal $\pi$ for $\max_{\pi}J_{\beta,f}(\pi,r^{*})$, we have
$\textrm{Term 1}\leq J_{\beta,f}(\pi^{*},r^{*})-J_{\beta,f}(\pi^{*},r_{t})\leq\sup_{x,y}\frac{\pi^{*}}{\pi_{t}}\mathbb{E}_{x\sim\rho,y\sim\pi_{t}}(r^{*}-r^{t}).$ (21)
Similarly, we can obtain its lower bound as $2R_{max}$. Then, we have
$\textrm{Term 1}\leq\sum_{t=1}^{T}\Big[\max_{\pi}\mathcal{J}_{\beta,f}(\pi,R(r_{t}))$ $-\max_{\pi}\mathcal{J}_{\beta,f}(\pi,R(r^{*}))\Big]+4R_{max}T.$ (22)
Bound term 2.
First, we utilize the preference generalized eluder coefficient to connect the prediction error to in-sample error.
$\textrm{Term 2}=\sum_{t=1}^{T}\mathbb{E}_{x\sim\rho,y\sim\pi^{t}(\cdot|x),y^{\prime}\sim\pi_{cal}}[f_{r_{t}}(x,y,y^{\prime})-f_{r^{*}}(x,y,y^{\prime})]$ (23)
$\leq\frac{\eta Td_{\mathrm{PGEC}}}{4}+4\sqrt{{d_{\mathrm{PGEC}}}T}+\frac{1}{\eta}\sum_{t=1}^{T}\mathbb{E}_{x\sim\rho,y\sim\widetilde{\pi}_{t}(\cdot|x),y^{\prime}\sim\pi_{cal}}[f_{r_{t}}(x,y,y^{\prime})-f_{r^{*}}(x,y,y^{\prime})]^{2},$ (24)
where the first equality uses the property of the reward class in Assumption C.2, and the inequality follows Definition C.1 with Cauchy–Schwarz inequality. Then, we bound the squared in-sample error as follows.
$\mathbb{E}_{x\sim\rho,\tau\sim\widetilde{\pi}_{t}}$ $[f_{r_{t}}(x,y,y^{\prime})-f_{r^{*}}(x,y,y^{\prime})]^{2}\leq\gamma\mathbb{E}_{x\sim\rho,\tau,\tau^{\prime}\sim\widetilde{\pi}_{t}}[f_{r_{t}}(x,y,y^{\prime})-f_{r^{*}}(x,y,y^{\prime})]^{2}$ (25)
$\leq\gamma(32R_{max}e^{4R_{max}})^{2}\mathbb{E}_{x\sim\rho,\tau,\tau^{\prime}\sim\widetilde{\pi}_{t}}[\sigma(f_{r_{t}}(x,y,y^{\prime}))-\sigma(f_{r^{*}}(x,y,y^{\prime}))]^{2}$ (26)
$\leq 8\gamma(32R_{max}e^{4R_{max}})^{2}\mathbb{E}_{x\sim\rho,\tau,\tau^{\prime}\sim\widetilde{\pi}_{t}(\cdot|x)}D_{H}^{2}(P_{r_{t}}(\cdot|\tau,\tau^{\prime})\|P_{r^{*}}(\cdot|\tau,\tau^{\prime})),$ (27)
where $\gamma=\sup_{x,y}\frac{\widetilde{\pi}_{t}}{\pi_{cal}}$, and the second inequality utilizes the Lemma C.8 in Xie et al. (2024), the third inequality uses $(x-y)^{2}<4(x+y)(\sqrt{x}-\sqrt{y})$. Refer to the Lemma C.6 in Xie et al. (2024), we have
$\sum_{i<t}\mathbb{E}_{x\sim\rho,\tau,\tau^{\prime}\sim\widetilde{\pi}_{t}(\cdot|x)}D_{H}^{2}(P_{r_{t}}(\cdot|\tau,\tau^{\prime})\|P_{r^{*}}(\cdot|\tau,\tau^{\prime}))\leq L_{BT}^{(t)}(r_{t})-L_{BT}^{(t)}(r^{*})+2\log(|\mathcal{R}|\delta^{-1})$ (28)
where $L_{BT}^{(t)}(r)=\sum_{i<t}\mathbb{E}_{y,y^{\prime}\sim\mathcal{D_{t}}}-\log\sigma(f_{r}(x,y,y^{\prime}))$ is the vanilla BT loss for reward modeling. Finally, the term 2 can be bounded by
$\textrm{Term 2}\leq 4\sqrt{{d_{\mathrm{PGEC}}}T}$ $+\frac{\eta Td_{\mathrm{PGEC}}}{4}+$
$\frac{8\gamma}{\eta}(32R_{max}e^{4R_{max}})^{2}(L_{BT}^{(t)}(r_{t})-L_{BT}^{(t)}(r^{*})+2T\log(|\mathcal{R}|\delta^{-1})).$ (29)
Bound the regret.
Since the $r_{t}$ is optimized by $L_{BT}^{(t)}(r_{t})-\sum_{i=1}^{T}\kappa\max_{\pi}\mathcal{J}_{\beta,f}(R(r_{t}),\pi)$, we have $r_{t}=\arg\min_{r\in\mathcal{R}}L_{BT}^{(t)}(r_{t})-\sum_{i=1}^{T}\kappa\max_{\pi}\mathcal{J}_{\beta,f}(R(r_{t}),\pi)$ Therefore, when $\eta=4\sqrt{\frac{2\gamma\log(T|\mathcal{R}|\delta^{-1})}{Td_{\mathrm{PGEC}}}}(32R_{max}e^{4R_{max}})$ and $\kappa=\frac{\eta}{4\gamma}(32R_{max}e^{4R_{max}})^{-2}$, the regret can be bounded by
Regret $\leq 4\sqrt{{d_{\mathrm{PGEC}}}T}+\sqrt{2^{3}\gamma d_{\mathrm{PGEC}}\log(|\mathcal{R}|\delta^{-1})}(32R_{max}e^{4R_{max}}T)+4R_{max}T.$ (30)
@@ Appendix C Regret Bound
Appendix D Experiments
D.1 Implementation Details

### v4

@@ Acknowledgement
We thank Froilan Choi and Shawn Im for their valuable suggestions on the paper, and thank all the ICLR 2026 PCs, SACs, ACs, and anonymous reviewers. This work is supported in part by the AFOSR Young Investigator Program under award number FA9550-23-1-0184, National Science Foundation under awards IIS-2237037 and IIS-2331669, Office of Naval Research under grant number N00014-23-1-2643, Schmidt Sciences Foundation, Open Philanthropy, Alfred P. Sloan Fellowship, and gifts from Google and Amazon.
@@
We have included all implementation details, hyperparameters, and training procedures in the paper and appendix. Our code and scripts for reproducing the experiments are publicly available through GitHub.
@@
This work studies reinforcement learning from human feedback (RLHF) using only publicly available or synthetic data, without new human subject collection. Here, by providing a rigorous theoretical framework with strong empirical evidence, we pursue a high standard of scientific excellence. We also take into account inclusiveness to make all our visualizations accessible to the unprivileged group of people by producing figures distinguished by light, shade, and marker. While RLHF has the potential to amplify biases or harmful behaviors if misused, our work is intended solely to advance safe and responsible research, and we encourage its application in alignment with ethical standards.
@@ Appendix A Related Works
Alignment (Hendrycks et al., 2021; Leike et al., 2018; Zhang et al., 2024b; 2025; Yeh et al., 2025) aims to ensure AI systems act in accordance with human values, preferences, and goals; and it has become a critical field in AI research. To steer language models to match human preferences, Reinforcement Learning from Human Feedback (RLHF) (Stiennon et al., 2020; Ouyang et al., 2022) achieves great success and has become the standard alignment pipeline. However, its computational complexity has motivated a family of Direct Preference Optimization (DPO) (Rafailov et al., 2024; Azar et al., 2024; Ethayarajh et al., 2024) that forgoes explicit reward modeling. Despite their efficiency, recent researchers (Tajwar et al., 2024; Xu et al., 2024; Xiong et al., 2024b) reemphasize the significance of online sampling.
@@
To address the computational overheads of passive exploration in RLHF, which samples trajectories just based on randomness, some existing attempts have been devoted to sample-efficient RL algorithms. Most of the works (Das et al., 2024; Ji et al., 2024; Dwaracherla et al., 2024; Mehta et al., 2023; Muldrew et al., 2024) adhere to the principle of optimism, proposing specialized prompt or response selection strategies to emphasize uncertain samples. While some research (Liu et al., 2024; Lou et al., 2024) propose uncertainty-aware reward models with epistemic neural networks or bootstrap ensembles, these methods introduce additional cost. Some research also addresses the sample efficiency with different theoretical foundations, such as information theory (Qi et al., 2025), preference-incentive exploration (Chen et al., 2025). Notably, several works (Zhang et al., 2024a; Xie et al., 2024; Cen et al., 2025) introduce different exploratory bonuses, which can implement optimism toward uncertainty without additional computations. However, they only focus on KL-divergence and their theoretical framework cannot result in real optimism as shown in Section 3.2.
@@
Beyond optimistic exploration, some research proposes fine-grained signals for RL learning. For instance, several studies propose segment-level (Kong et al., 2025; Yin et al., 2025) or token-level (Zeng et al., 2024; Li et al., 2024) reward functions for alignment or text control. Notably, for reasoning tasks, the process reward model (Lightman et al., 2024; Li and Li, 2025; Yuan et al., 2024), which provides step-wise feedback for solutions, has shown promising effectiveness. On the other hand, recent research (Cheng et al., 2025; Zheng et al., 2025; Wu et al., 2025) on LLM reasoning reveals that high-entropy tokens guide the model toward diverse reasoning paths. Training with only high-entropy tokens is more beneficial for reasoning performance (Wang et al., 2025). While our approach is highly extensible, we believe the orthogonal methods can be further incorporated with our general exploratory bonus.
@@ Appendix B Optimism Failure of previous works
B.1 Optimism failure under KL-divergence
We start by proving how the existing exploratory bonus term under KL-divergence instantiation fails to achieve optimistic exploration.
Lemma 3.1
Proof  First, the inner maximization of the bonus term admits a closed-form solution, $\pi^{*}(y|x)=\pi_{\textrm{ref}}(y|x)e^{\frac{r(x,y)}{\beta}}/Z(x)$ where $Z(x)=\mathbb{E}_{y\sim\pi_{\textrm{ref}}(\cdot|x)}e^{\frac{r(x,y)}{\beta}}$ is a normalization term. Substituting this solution for the bi-level objective of $r_{2}$ reduces it to a single-level form:
As shown in Rafailov et al. (2024), the log-ratio $\beta\log\pi_{\theta}(y|x)-\beta\log\pi_{\textrm{ref}}(y|x)$ represents the same class of the original reward function $r$ through Eq. 4, thus all $r$ in the reward modeling objectives can be reparameterized by the log-ratio. Plugging this into Eq. 12 yields
Since the second term equals 0, the reparameterized Eq. 12 is exactly the vanilla DPO loss, that is, the reparameterized training objective of $r_{1}$. Thus, the exploratory bonus in the reward training objective does not affect the final policy set. $\square$
B.2 Extension beyond $\alpha$-divergence
@@
[…]
Proof  By Lemma 3.2, we can reparameterize the bonus term used for optimistic reward modeling into Eq. 6. Denote $h(u)=uf^{\prime}(u)-f(u)$. For a fixed prompt $x$, the training step can then be written as the following constrained optimization problem:
@@
where $\mu,\eta$ are the dual variables. Then we utilize the Karush-Kuhn-Tucker (KKT) conditions for the given optimization problem. The complementary slackness gives that $\forall y,\eta(y)\pi(y|x)=0$. The stationary condition requires
@@
Let $S_{y}=\{y|\pi(y|x)>0\}$. For all $y\in S_{y}$, complementary slackness implies $\eta(y)=0$. Since $h^{\prime}(u)=uf^{\prime\prime}(u)$ is a monotone function, we can obtain $\forall y\in S_{y},\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)}$ is a constant. Then applying the normalization constraint, $\mathbb{E}_{y\sim\pi_{\textrm{ref}}(\cdot|x)}\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)}=1$. Hence, we conclude that the unique interior optimum is $\pi^{*}(y|x)=\pi_{\textrm{ref}}(y|x)$. $\square$
@@
B.3 Prior adaptions & approximations cannot generalize
Although the theoretical justification for the previously proposed exploratory bonus does not hold, its empirically implemented loss remains effective due to various adaptations and approximations. In this subsection, however, we show that these adaptations and approximations cannot be extended beyond the KL-divergence class.
Zhang et al. (2024a) modify the formulation of $\mathcal{J}_{\beta,f}(\pi,r)$ in the optimistic exploratory bonus $\kappa\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r)$ as
@@ B.3 Prior adaptions & approximations cannot generalize
which adds a bias in the reward expectation term. Under KL-divergence, the original $\mathcal{J}_{\beta,f}(\pi,r)$ will be zero after re-parameterization as shown in Lemma 3.1, thus the sole reparameterized bias term will remain as $\mathbb{E}_{y^{\prime}\sim\pi_{\textrm{ref}}}\log\pi(y^{\prime}|x)$. Since $\mathcal{J}_{\beta,f}(\pi,r)$ cannot be reparameterized to zero except KL-divergence, this adaptation cannot generalize then.
In the derivations of Cen et al. (2025), an idealized calibration distribution $\pi_{cal}$ is assumed to satisfy $\mathbb{E}_{y\sim\pi_{cal}}r(x,y)=0$. Since $\pi_{cal}$ is not directly accessible in practice, the method substitutes rejected responses to approximate expectations under $\mathbb{E}_{\pi_{cal}}$. These rejected samples, however, do not satisfy the defining property of $\pi_{cal}$, making the approximation theoretically inconsistent.
The theoretical framework of Xie et al. (2024) is based on Implicit Q-Approximation (refer to Lemma C.3 in the original paper) as follows,
$\beta\frac{\log\pi(y|x)}{\log\pi_{\textrm{ref}}(y|x)}=r(x,y)-V^{*}(x)$ (18)
where $V(x)^{*}$ is the KL-regularized value function. Because this relation depends critically on the logarithmic structure of the KL divergence, the framework cannot be extended to more general divergence families either.
@@
In contrast, our general exploratory bonus integrates seamlessly with iterative online RLHF algorithms and extends naturally to the entire $\alpha$-divergence family. As discussed in §4, all these heuristic bonus terms can be encompassed by our unified theoretical framework.
B.4 Equivalence between a sophisticated objective and a simple one.
Lemma 4.1
Denote two objectives as $h(u(x,y))=\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}u(x,y)f^{\prime}(u(x,y))-f(u(x,y))$ and $\hat{h}(u(x,y))=\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}\frac{u(x,y)}{Z(x)}f^{\prime}(u)-f(\frac{u(x,y)}{Z(x)})$ where $u(x,y)$ is a function with $\pi(y|x)$ and $\pi_{\textrm{ref}}(y|x)$. If the ratio
$\frac{f^{\prime}(u(x,y))+u(x,y)f^{\prime\prime}(u(x,y))-f^{\prime}(\frac{u(x,y)}{Z(x)})}{Z(x)u(x,y)f^{\prime\prime}(u(x,y))}=\Lambda(x)$ (19)
is independent of $y$ and $\Lambda(x)>0$, then minimizing the two objectives, $\min_{\pi}-h(u(x,y))$ and $\min_{\pi}-\hat{h}(u(x,y))$, yields the same class of optimal policies.
Proof  Similar to Lemma 4.1, for a fixed $x$, we can write a similar formulation of the Lagrange multipliers as follows,
$\mathcal{L}=\mathbb{E}_{y\sim\pi_{\textrm{ref}}}h(u(x,y))-\mu_{1}(\sum_{y}\pi(y|x)-1)-\sum_{y}\eta_{1}\pi(y|x),$ (20)
where $\mu,\eta$ are the dual variables. Then, we utilize the KKT conditions formulation for the given optimization problem. Similar to Lemma 4.1, when $\pi(x,y)>0$, we obtain
$\frac{\partial h}{\partial\pi(y|x)}$ $=u(x,y)f^{\prime\prime}(u(x,y))\cdot\pi_{\textrm{ref}}(y|x)\cdot\frac{\partial u(x,y)}{\partial\pi(y|x)}=\mu_{1}.$ (21)
Similarly, we can obtain the KKT conditions for $\hat{h}(u(x,y))$ as follows,
$\frac{\partial\hat{h}}{\partial\pi(y|x)}=\frac{1}{Z(x)}(f^{\prime}(u(x,y))+u(x,y)f^{\prime\prime}(u(x,y))-f^{\prime}(\frac{u(x,y)}{Z(x)}))\cdot\pi_{\textrm{ref}}(y|x)\cdot\frac{\partial u(x,y)}{\partial\pi(y|x)}=\mu_{2},$ (22)
where $\mu_{2}$ is another dual variable. With the condition of Eq. 20, these two partial derivatives in Eq. 21 and Eq. 22 are equivalent when $\mu_{2}=\mu_{1}\Lambda(x)$. Hence, every policy that satisfies the stationary condition for $h$ also satisfies it for $\hat{h}$. Since $\Lambda(x)>0$, the second-order derivative $\frac{\partial^{2}h}{\partial^{2}\pi(y|x)}$ and $\frac{\partial^{2}\hat{h}}{\partial^{2}\pi(y|x)}$ have the same sign, which indicates they share the same local minima. Hence, minimizing the two objectives $\min_{\pi}-h(u)$ and $\min_{\pi}-\hat{h}(u)$ induces the same class of policies.
@@
B.5 GEB enables optimistic exploration
In this subsection, we prove how GEB meets the optimism principle specified by Definition 3.1.
Theorem 4.2
Consider an $\alpha$-divergence $f$ with $0\leq\alpha\leq 1$, and the exploratory bonus $\mathcal{L}_{\text{bonus}}=\beta\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}\Big[u(x,y)f^{\prime}(u(x,y))-f(u(x,y))\Big]$, where $u(x,y)$ is a function dependent on $\pi(y|x)$ and $\pi_{\textrm{ref}}(y|x)$. For any $(x,y)$, if $\frac{\partial u}{\partial\pi}+\pi_{\textrm{ref}}\frac{\partial^{2}u}{\partial\pi\partial\pi_{\textrm{ref}}}+\frac{(\alpha-1)\pi_{\textrm{ref}}}{u}\frac{\partial u}{\partial\pi}\frac{\partial u}{\partial\pi_{\textrm{ref}}}<0$ and $u(x,y)>\alpha$, the optimism condition in Definition 3.1 is satisfied; that is, $\frac{\partial^{2}\mathcal{L}_{\text{bonus}}}{\partial\pi\partial\pi_{\textrm{ref}}}\leq 0$.
$\mathcal{L}_{\text{bonus}}=\beta\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}(\cdot|x)}\Big[u(x,y)f^{\prime}(u(x,y))-f(u(x,y))\Big].$ (23)
For $\alpha$-divergences, the conditions in Lemma 4.1 are satisfied. Since we have $f^{\prime}(u)+uf^{\prime\prime}(u)-f^{\prime}(\frac{u}{Z})=u^{\alpha-1}(Z^{1-\alpha}-\alpha)/(1-\alpha)$ and $uf^{\prime\prime}(u)=u^{\alpha-1}$, the fraction $\Lambda(x)=(Z^{1-\alpha}-\alpha)/Z(1-\alpha)$ in Lemma 4.1 is independent of $y$. Since $u>\alpha$, $Z_{R}=\mathbb{E}_{y\sim\pi_{\textrm{ref}}(\cdot|x)}u>0$, thus $\Lambda(x)>0$ is also satisfied. Finally, the mixed second-order derivative of Eq. 11 is computed as
which achieves the optimism defined in Definition 3.1. $\square$
Appendix C Experiments
C.1 Implementation Details